A two- digit is such that the sum of the digits is 11. When the number with the same digits reversed is subtracted from this number, the difference is 9. What is the number?
Answer by lwsshak3(3703)

A two- digit is such that the sum of the digits is 11. When the number with the same digits reversed is subtracted from this number, the difference is 9. What is the number?
**
let u=units digit
let t=tens digit
u+t=11
t=11-u
original number=10t+u
reversed number=10u+t
10t+u-10u-t=9
10(11-u)+u-10u-(11-u)=9
110-10u+u-10u-11+u=9
18u=90
u=5
t=11-u=6
ans:
original number=65